OpenStudy (ell):

Help please!! :) Solve the equation by factoring: 7r^2-14r=-7

OpenStudy (anonymous):

7r² - 14r + 7 = 0 7(r² - 2r + 1) = 0 7(r-1)(r-1) = 0 FACTORED r = 1

OpenStudy (ell):

thanx!:) but I thought there was supposed to be 2 answers....

OpenStudy (anonymous):

YW:) If there were two variables in the question, then you could solve for both, but here, you can only find r..

OpenStudy (ell):

oh! I see, haha(: thanks! Maybe you can help with this one?? 5r^2-44+120=-30+11